OpenStudy (anonymous):

the expression radical -50 / radical 2

OpenStudy (anonymous):

\[\sqrt{-50} / \sqrt{2}\]

OpenStudy (mathmale):

Please include all of the instructions. What's your goal?

OpenStudy (anonymous):

the expression \[\sqrt{-50} / \sqrt{2}\] is equivalention to a 5i b -5 c 5 d-5i @mathmale

OpenStudy (mathmale):

Much depends upon your understanding that Sqrt(-1) = i, the imaginary operator. Have you seen this before?